### v1

# Dark energy as a battery for magnetic field generation in plasmas

Abstract
It is shown that in the spacetime dominated by a cosmological constant, in the far region of a Schwarzschild-de Sitter black hole, a seed magnetic field can be generated in an ambient plasma (in a state of no magnetic field) by a general-relativistic battery, which depends on the interaction of spacetime curvature with inhomogeneous plasma thermodynamics. Thus, at large distances, dark energy becomes the only gravitational source for magnetic field generation. This allows a mechanism that make dark energy manifest through its conversion to cosmic magnetic fields.
When the primary source of magnetic fields - the electrical current - is not explicitly evident in a system, one must resort to indirect sources with free energy that could generate magnetic energy. There is a large body of literature on the so-called dynamo mechanisms (see, for instance, Refs. KR ; MGM ; MGM2 ; MS ) that could amplify a magnetic field from a non-zero seed. Therefore, looking for processes/systems that could generate a seed magnetic field has been a serious quest in plasma physics. Let us call such seed creation systems, collectively, as Batteries; the most famous of these is the very original Biermann Battery (BB) biermann ; kulr . The seed magnetic field in BB originates in the inhomogeneous thermodynamics (of a special type) that harnesses the free energy in density and temperature gradients.
In principle, the Biermann battery can exist for plasmas at all energy scales. However, it is neither unique nor always dominant; when special and/or general relativistic effects are considered, new batteries appear yoshida ; asenjoMahajan ; M2 ; asenjoMahajan2 and also may dominate. These batteries may originate from different free energy sources. For instance, the source of magnetic energy in Refs. yoshida ; M2 lies in the free energy in the velocity shear. However, in Refs. asenjoMahajan ; asenjoMahajan2 , the possibility of using the spacetime curvature as the source for magnetic field generation is explored. These relativistic batteries are likely more effective in “extreme conditions”.
In this work, we explore precisely one of these systems, the plasma dynamics in the background of a Schwarzschild-de Sitter black hole. We plan to demonstrate the working of the general relativistic battery asenjoMahajan driven by the interaction of this spacetime curvature with plasma thermodynamics. In particular, we will focus on the vortical plasma dynamics very far from the black hole - in the region where the spacetime curvature is controlled entirely by the cosmological constant (dark energy), as the curvature produced by the black hole mass is negligible. To our knowledge, this is the first calculation connecting dark energy to magnetic field generation (seed).
The spacetime curvature of the Schwarzschild-de Sitter black hole rindler is described by the metric $g_{\mu\nu}$; its non-vanishing components (in spherical coordinates) are $g_{tt}=-\alpha^{2}$, $g_{rr}=\alpha^{-2}$, $g_{\theta\theta}=r^{2}$, and $g_{\phi\phi}=r^{2}\sin^{2}\theta$, with
$\alpha=\left(1-\frac{2M}{r}-\frac{\Lambda r^{2}}{3}\right)^{1/2}\,,$ (1)
where $M$ is the black hole mass, and $\Lambda$ is the positive cosmological constant of a de Sitter Universe. This metric allows the realization of a black hole in an expanding Universe. In this spacetime, several plasma propagation modes have been studied Hossain ; sharif ; Atiqur .
In this work, we study the dynamics of a plasma on the spacetime background given by the previous metric. In general, the covariant dynamics of a non-gravitating plasma fluid (the intrinsic gravity of the plasma is neglected) in a curved spacetime background can be described in terms of a unified field $M^{\mu\nu}$ M1-M3 ; asenjoMahajan . The dynamics can be put in the unified form
$U_{\nu}M^{\mu\nu}=\frac{T}{q}\nabla^{\mu}\sigma\,,$ (2)
where $M^{\mu\nu}=F^{\mu\nu}+(m/q)S^{\mu\nu}$ is a fully antisymmetric tensor comprising the electromagnetic tensor $F^{\mu\nu}$. The plasma, composed by particles with mass $m$, and charge $q$, is characterized by the tensor $S^{\mu\nu}=\nabla^{\mu}(fU^{\nu})-\nabla^{\nu}(fU^{\mu})$, with four-velocity $U^{\mu}$, and temperature $T$. Besides, $\sigma$ ($f$) are the entropy (enthalpy) per mass densities M1-M3 . Lastly, $\nabla_{\mu}$ is the covariant derivative defined for a general metric $g_{\mu\nu}$. The dynamics is complete when complemented by the generally covariant Maxwell equations.
From the basic Eq. (2), the expression for magnetic field generation (in a spherically symmetric static spacetime) can be readily derived by following the procedure detailed in Ref. asenjoMahajan . One affects a $3+1$ decomposition of spacetime, projecting all physical quantities onto timelike and spacelike hypersurfaces; the spacetime metric is decomposed as $g_{\mu\nu}=\gamma_{\mu\nu}-n_{\mu}n_{\nu}$, where $n_{\mu}$ is a normalized timelike vector that satisfies $n_{\mu}n^{\mu}=-1$, and $n^{\mu}\gamma_{\mu\nu}=0$. Similarly, $\gamma_{\mu\nu}$ is the 3-metric of the spacelike hypersurfaces of metric $g_{\mu\nu}$. For the spherically symmetric static spacetime of the Schwarzschild-de Sitter black hole, the metric acquires the form $g_{\mu\nu}=(-\alpha^{2},\gamma_{ij})$ yielding $n_{\mu}=(\alpha,0,0,0)$, $\gamma_{rr}=\alpha^{-2}$, $\gamma_{\theta\theta}=r^{2}$, and $\gamma_{\phi\phi}=r^{2}\sin^{2}\theta.$ By combining the spacelike projection of Eq. (2) with the spacelike projection of the identity $\nabla_{\nu}M^{*\mu\nu}=0$ (where $M^{*\mu\nu}$ is the dual of $M^{\mu\nu}$), the generalized plasma vorticity equation is found to be asenjoMahajan
$\frac{\partial{\bf\Omega}}{\partial t}-\nabla\times\left({\bf v}\times{\bf\Omega}\right)={\bf\Xi}_{B}+{\bf\Xi}_{R}\,,$ (3)
where the generalized plasma vorticity ${\bf\Omega}$ is the vectorial part of the spacelike tensor $\Omega^{\mu}=(1/2)n_{\rho}\epsilon^{\rho\mu\sigma\tau}M_{\sigma\tau}$, with the totally antisymmetric tensor $\epsilon^{\rho\mu\sigma\tau}$ (notice that $n_{\mu}\Omega^{\mu}=0$). Also, ${\bf v}$ is the plasma fluid velocity, being the vectorial part of the four-velocity $v^{\alpha}=(1/\Gamma){\gamma^{\alpha}}_{\mu}U^{\mu}$. Here, $\Gamma=(1/\alpha)n_{\mu}U^{\mu}$ is the corresponding Lorentz factor of the fluid velocity, such that the four-velocity can be written as $U^{\mu}=-\alpha\Gamma n^{\mu}+\Gamma{\gamma^{\mu}}_{\nu}v^{\nu}$. Thus, $\Gamma=(\alpha^{2}-\gamma_{\mu\nu}v^{\mu}v^{\nu})^{-1/2}$. Besides, the $\nabla$ operator corresponds to the spacelike projection of the covariant derivative. Furthermore, the vectorial components of the generalized plasma vorticity can be written as
${\bf\Omega}={\bf B}+\frac{m}{q}\nabla\times\left(f\Gamma{\bf v}\right)\,,$ (4)
where ${\bf B}$ is the vectorial part of spacelike magnetic field tensor, defined as
$B^{\mu}=\frac{1}{2}n_{\rho}\epsilon^{\rho\mu\sigma\tau}F_{\sigma\tau}\,.$ (5)
Also, the term ×$\nabla\times\left(f\Gamma{\bf v}\right)$ in Eq. (4) corresponds to the vectorial part of the fluid vorticity.
The two terms on the right-hand side of Eq. (3) are the batteries of the theory. This equation predicts that these batteries can generate generalized vorticity. The first one is the general relativistic-corrected Biermann battery asenjoMahajan
${\bf\Xi}_{B}=-\frac{1}{q\Gamma}\nabla T\times\nabla\sigma\,,$ (6)
that depends only on the non-parallel spatial variations of the plasma temperature and entropy. Both quantities tend to have parallel gradients in most simple cases, and the Biermann battery vanishes. The second battery
${\bf\Xi}_{R}=\frac{T}{q\Gamma^{2}}\nabla\Gamma\times\nabla\sigma\,,,$ (7)
is the general relativistic drive asenjoMahajan , the prime objective of this paper. Notice that this battery appears by the interplay between relativistic kinematical effects, spacetime curvature, and the thermodynamical properties of the plasma; the spacetime curvature is also embedded in the definition of $\Gamma$.
In this unified formalism, the electromagnetic field and the fluid thermal-vortical field appear together, and what is being generated via Eq. (3) is the generalized vorticity (or the generalized magnetic field). In the conventional non-relativistic theories like the original BB, the thermal-vortical part is neglected, and the generalized vorticity is reduced to the magnetic field M1-M3 .
The generalized vorticity equation (3) and the importance of the general relativistic drive (7) has been explored in several astrophysical scenarios, such as plasmas around Schwarzschild black holes asenjoMahajan ; kana , Reissner-Nördstrom and Kerr black holes asenjoMahajan2 ; Bhattacharjee1 ; Bhattacharjee2 ; Bhattacharjee3 ; Bhattacharjee4 . In all those cases, the mass of the black hole is the source of the space-time curvature; the magnetic field generation is, thereby, a local phenomenon. Here, we make a fundamental departure in exploring the effect of the pervading dark energy through the cosmological constant.
To calculate the effect of the cosmological constant on the generation of the magnetic field seed, and in the same spirit as the earlier general relativistic calculations, we assume that the general relativistic drive (7) is more relevant than the Biermann battery (6). Without loss of generality, we consider the metric (1), evaluated at the black hole’s equatorial plane ($\theta=\pi/2$), to study the plasma dynamics. Far from the black hole, let us assume that the fluid has only radial velocity, ${\bf v}=v^{r}\hat{e}_{r}$, and $v^{r}=v^{r}(r)$. For this choice, the fluid vorticity vanishes identically, and the generalized vorticity (4) reduces to the magnetic field, ${\bf\Omega}={\bf B}$.
Due to the chosen symmetry of our system, let us consider the generation of a magnetic field with a polar component in the equatorial plane. The components of the magnetic field can be explicitly evaluated using Eq. (5). This allows us to obtain that the orthogonal component to the equatorial plane is ${B}^{\theta}=\alpha B^{\theta}_{\mbox{\footnotesize{flat}}}$, where $B^{\theta}_{\mbox{\footnotesize{flat}}}$ is the polar magnetic field (in the equatorial plane) in flat spherical coordinates. Thereby, the (only non-vanishing) polar component of Eq. (3) can be written as
$\frac{\partial{B}^{\theta}_{\mbox{\footnotesize{flat}}}}{\partial t}+\frac{1}{r^{2}}\frac{\partial}{\partial r}\left({r^{2}}v^{r}B^{\theta}_{\mbox{\footnotesize{flat}}}\right)=\frac{{\Xi}_{B}^{\theta}}{\alpha}+\frac{{\Xi}_{R}^{\theta}}{\alpha}\,.$ (8)
Now, let us first consider the simplest case where the Biermann battery (6) vanishes. This occurs when $\nabla T\parallel\nabla\sigma$. In such cases, the only available battery is the one given by (7). To evaluate this battery, let us consider an initial state with no plasma nor magnetic field. Thus, we can consider an initial flow of particle constituents moving radially, with velocity $v^{r}=U^{r}/\Gamma=\dot{r}/\Gamma=\sqrt{E^{2}-\alpha^{2}}/\Gamma$ jak , where $E$ is a dimensionless constant related to the energy density of the particle, determined at the point $r_{0}$ when the particle is at rest, $E=\alpha(r_{0})$]. Besides, the Lorentz factor is $\Gamma={E}/{\alpha^{2}}$ jak . With this, we can calculate the relativistic drive (7), which is non-zero only if angular gradients of the entropy are allowed. The polar component of the relativistic drive, then, is
${\Xi}_{R}^{\theta}=\frac{T\alpha^{5}}{qEr}\frac{\partial}{\partial r}\left(\frac{1}{\alpha}\right)\frac{\partial\sigma}{\partial\phi}\,,$ (9)
which, very far from the black hole, becomes
$\left.\frac{{\Xi}_{R}^{\theta}}{\alpha}\right|_{r\ggg 2M}\approx\frac{T\,\Lambda}{3qE}\frac{\partial\sigma}{\partial\phi}\,,$ (10)
scaling linearly with the cosmological constant $\Lambda$. Notice that, for this battery not to diverge at large distances, the temperature $T$ and entropy $\sigma$ must remain bounded.
In contrast, one may wonder under what conditions can the Biermann battery compete with the cosmological constant induced relativistic drive (10). If the temperature and entropy gradients are not parallel, the polar component of the Biermann battery (6), ${\Xi}_{B}^{\theta}=-({\alpha^{3}}/{qEr}){\partial_{r}T}{\partial_{\phi}\sigma}$, becomes (far away from the black hole)
$\left.\frac{|{\bf\Xi}_{B}^{\theta}|}{\alpha}\right|_{r\ggg 2M}\approx\frac{\Lambda}{3qE}r\frac{\partial T}{\partial r}\frac{\partial\sigma}{\partial\phi}\,.$ (11)
implying
$|{\bf\Xi}_{B}^{\theta}|\ll|{\bf\Xi}_{R}^{\theta}|\,,$ (12)
as long as $r\partial_{r}T\ll T$. The Biermann battery will be negligible at large distances, and the dark energy-dominated part of the general relativistic drive will dominate the magnetic field generation. Let us name this drive as the Dark Energy Battery (DEB).
Let us go back to Eq. (8) to estimate the magnetic field generated by DEB starting from a state of no magnetic field. For a general time evolution, the calculation will involve solving complicated nonlinear equations. In fact, a self-consistent solution can be affected by solving Eq. (8) simultaneously with equations evolving the entropy evolution. This work will necessarily involve numerical solutions of the system and will be deferred to a later publication. This paper, devoted to the first recognition of a new pervading drive, will do the local estimate of the linear phase of the seed generation.
In this regime, because battery (10), only the cosmological constant is responsible for generating magnetic fields. In the very-far limit ($r\ggg 2M)$, for times measured from an arbitrary $t=0$, the rate of magnetic field generation is
$\frac{\partial}{\partial t}{B^{\theta}_{\mbox{\footnotesize{flat}}}(r\ggg 2M)}\approx\frac{T\Lambda}{3qE}\frac{\partial\sigma}{\partial\phi}\,,$ (13)
which remains well-defined as the thermodynamical quantities remain finite. Notice that the only gravitational (source of spacetime curvature) entity in Eq. (13) is the cosmological constant $\Lambda$ representing dark energy.
Eq. (13) gives the non-vanishing initial condition required to generate a magnetic field. Thus, the magnetic field grows linearly with time until a time $\tau$, when the nonlinearities of the vortical dynamics [second term at the left-hand side of Eq. (8)] become relevant. This time can be estimated to be $\tau\sim|L/v^{r}|$, where $L$ is the length scale of radial variations of the magnetic field asenjoMahajan . The condition for the radial variations of $B^{\theta}_{\mbox{\footnotesize{flat}}}$ to be relevant in this spacetime is $1/L\sim|\partial_{r}B^{\theta}_{\mbox{\footnotesize{flat}}}/B^{\theta}_{\mbox{\footnotesize{flat}}}|\sim|\partial_{r}\alpha/\alpha|$. For the metric (1), very far from the black hole, we find that $L\sim r$, implying that magnetic fields present almost no variations at infinity. With this, at large distances, the time scale (far from the black hole), then, goes as
$\left.{\tau}\right|_{r\ggg 2M}\sim\frac{E}{r^{2}}\left(\frac{3}{\Lambda}\right)^{3/2}\,.$ (14)
At time $\left.{\tau}\right|_{r\ggg 2M}$, the magnetic field will grow to its maximum value (permitted in this model) given by
$B^{\theta}_{\mbox{\footnotesize{max}}}(r\ggg 2M)\approx\frac{T}{q\,r^{2}}\sqrt{\frac{3}{\Lambda}}\frac{\partial\sigma}{\partial\phi}\,,$ (15)
leading to the expected spherical $1/r^{2}$ behavior. This implies that the generated covariant polar (angular) magnetic field $B_{\theta}=g_{\theta\theta}B^{\theta}$, very far, depends only on thermodynamics and dark energy
$B_{\theta\mbox{\footnotesize{max}}}(r\ggg 2M)\approx\frac{T}{q}\sqrt{\frac{3}{\Lambda}}\frac{\partial\sigma}{\partial\phi}\,.$ (16)
In this model calculation, we have shown that the cosmological constant (pervading dark energy) can trigger the generation of a magnetic field (15) in a plasma present in the outer reaches of a black hole, as long as the plasma entropy density has local azimuthal variations. Like the conventional BB battery, this new battery is powered by the interaction of gradient-free energy with the space-time curvature induced by the dark energy. The linear growth stage (from an initial state of no magnetic field) lasts for a time $\tau$ during which the field reaches a maximum value (15). About and beyond the time $\tau$, a much more elaborate calculation is needed - we plan to do it in a later publication.
The $B^{\theta}_{\mbox{\footnotesize{max}}}$ generated by this mechanism (in the region very far from the black hole) is expected to be rather small. Although the value of the cosmological constant $\Lambda$ is very small, the azimuthal variations of entropy (and, therefore, of temperature) tend to be also small. Besides, the field ends up decaying as $1/r^{2}$ at infinity.
It is to be emphasized that, in the spacetime considered here (dominated by dark energy), the general relativistic drive DEB (10) dominates the Biermann battery and is the primary cause of the finite magnetic field seed (13), proportional to ${\Lambda}$. This is the principal result of this work, where we explored a unique new mechanism where dark energy can be converted into magnetic energy. This calculation establishes a basis (in principle) for detecting the signatures of dark energy through electromagnetic energy. By the same token, if there is dark energy, we will always have a source for cosmic magnetic fields.
The very idealized simple estimates made in this paper need to be supplemented by a solution of the nonlinear Eq. (3) [or Eq. (8)] to get a better idea about the saturation of the magnetic field. This will be allowed by fully coupling the dynamics to the Maxwell equations. Such a system will have to be solved numerically and will be the subject of a more detailed paper.
The above simple analysis, however, has provided ample support for the mechanism that may convert the dark energy into well-known and well-understood magnetic energy.
Acknowledgments
We are grateful to ANID-Chile for Doctoral Scholarship No. 21220616 (NVS), and Fondecyt grants No. 1240281 (PSM) and No. 1230094 (FAA) that partially supported this work. SMM’s research is supported by US DOE Grants DE-FG02-04ER54742 and DE-AC02-09CH11466.